OpenStudy (anonymous):

John and Barry are pulling on a fence post with ropes while their friend Pat uses a chain-saw to sheer it at ground-level. John applies a force of 75.0N [40*] on one rope while Barry applies a force of 85.0N [340*] on the other rope. What is the net force on the post? Show your work.

OpenStudy (anonymous):

x component of forces: 75 cos (40) + 85 cos (20)=137.3 y component 75 sin (40) -85 sin (20) = -31.7 magnitude = square root of sum of the components squared, 140.9 angle is that whose tangent is =31.7/137.3, -13o or 347o.

OpenStudy (anonymous):

@douglaswinslowcooper would you mind drawing a picture? I kind of get it but kind of don't...

OpenStudy (anonymous):

cos(340)=cos(20) sin(340)=-sin(20)